Verify: cos3x=cos³x-3sin²xcosx
1
Expert's answer
2019-12-16T09:59:21-0500

cos3x=cos(2x+x)=cox2x*cosx-sin2x*sinx=(1-2sin2x)*cosx-2sinx*cosx*sinx=

=cosx-2sin2x*cosx-2sin2x*cosx=cosx*(1-4*(1-cos2x))=cosx*(1-4+4cos2x)=4cos3x-3cosx;

cos3x=cos3x-3sin2x*cosx

4cos3x-3cosx=cos3x-3sin2x*cosx;

4cos3x-3cosx-cos3x+3sin2x*cosx=0;

3cos3x+3sin2x*cosx-3cosx=0;

cosx(3cos2x+3sin2x-3)=0;

cosx*(3*(cos2x+sin2x)-3)=0;

cosx*(3*1-3)=0;

cosx*0=0, which is true, hence the initial formula was verified.